Tax. Determine the SWL.
D: P = 440 - 2Q
S: P = 40 + 2Q
Tax: P = 80
1
Expert's answer
2017-07-11T03:27:07-0400
440-2Q=40+2Q
4Q=400
Q=100
SWL=P(tax)*Q=80*100=8000